# Mixture-of-Subspaces in Low-Rank Adaptation

### v1

@@ 4.1 Commonsense Reasoning
We fine-tune LLaMA-3 8B instruction version model (AI@Meta, 2024) for the commonsense reasoning question answering tasks. We first train the model using 170k training samples (Hu et al., 2023), and then test the fine-tuned model on 8 commonsense reasoning question answering benchmarks (refer to Appendix A.1 for details). The 170k training set is the mixture of the training sets of these benchmarks. Besides LoRA (Hu et al., 2022), we also compare MoSLoRA with various baselines, including: 1) LoKr (Yeh et al., 2023) which employs Kronecker products for matrix decomposition of $\mathbf{A}\mathbf{B}$; 2) LoHa (Yeh et al., 2023) which decomposes the vanilla LoRA into the Hadamard product of two LoRA branches; 3) FLoRA (Si et al., 2024) which introduces an extra core based on Tucker decomposition to maintain the consistent topological structure with the original space  [^3] Please refer to Appendix D for the discussion of differences. 4) AdaLoRA (Zhang et al., 2023) which parameterizes the incremental updates of the pre-trained weight matrices in the form of singular value decomposition; and 5) DoRA (Liu et al., 2024) which decomposes the pretrained weight into its magnitude and directional components and fine-tunes both of them.
@@ 4.2 Visual Instruction Tuning
There are two stages in training LLaVA: 1) pretrain a two-layer MLP to project visual features to language space, and 2) optimize LLM and visual encoder (optional) for visual instruction tuning. In this paper, we employ the pretrained projector provided in XTuner (Contributors, 2023b), and then conduct visual instruction tuning on the LLM backbone and visual encoder, simultaneously. For the LLM backbones, we adapt the LLaMA3 8B (AI@Meta, 2024) and InternLM2 7B (Cai et al., 2024) using the off-the-shelf projecters 444pretrained projecters. For the visual encoder, we employ the ViT 555openai/clip-vit-large-patch14-336 (Dosovitskiy et al., 2021) large version. Due to limited resources, we fintune both the LLM backbone and visual encoder via LoRA/MoSLoRA on the 665K instruction-following data (Liu et al., 2023a), rather than optimize all the parameters in LLMs. For InternLM2, we employ the 4-bit QLoRA (Dettmers et al., 2023) and corresponding QMoSLoRA (QLoRA+MoSLoRA). Based on the results in Table 3, we initialize the mixer as the orthogonal matrix and Kaiming uniform distribution, separately. For specific hyperparameters, please refer to the Appendix B.2. It takes around 20 hours to fine-tune using 4 Nvidia A100 80G GPUs.
@@
After visual instruction tuning, we evaluate the trained model on 9 popular benchmarks, including MMBench EN/CN (Liu et al., 2023b), SEED Bench (Li et al., 2023), AI2D (Kembhavi et al., 2016), SciQA (Lu et al., 2022), TextVQA (Singh et al., 2019), MathVista testmini (Lu et al., 2023), MM-Vet (Yu et al., 2023), and MME (Fu et al., 2023). All the evaluations are done using the VLMEvalKit (Contributors, 2023a). Please refer to Appendix A.2 for the details of the dataset and the reported metrics. Specifically, we scale the MME scores to 100 to calculate the average score.
@@ 4.3 Subject-driven Generation
Figure 6 shows one case of a dog from the DreamBooth dataset (Ruiz et al., 2023). We finetune the SDXL666stable-diffusion-xl-base-1.0 model (Podell et al., 2023) via LoRA and MoSLoRA. In MoSLoRA, the mixer is initialized as an orthogonal matrix. During finetuning, the learning rate is 1e-4, and the batch size is 4. We train the model for 500 steps, which costs around 16 minutes using 1 80G A100 GPU. During generation, we infer 50 steps for the given prompts. Compared to vanilla LoRA, we can find that our proposed MoSLoRA captures more details of the subject and better conforms to the given prompt. Specifically, MoSLoRA learns more details about the given cat, including the color of the hairs around the neck and the shape of the paw. Meanwhile, the images from MoSLoRA are more consistent with the given prompts, such as the wet (thus clumped) hair and the floating gesture (spread hands).
@@
We also perform human evaluation on the generated images. First, we choose four subjects (i.e. cat, dog, grey sloth plushie, and can) from the DreamBooth dataset (Ruiz et al., 2023) and fine-tune the SDXL model, respectively. Then, we randomly select 8 prompts to generate the corresponding images. After that, 15 human experts are asked to independently score win/tie/loss for the paired images from LoRA and MoSLoRA. During evaluation, we shuffle these pairs and keep that these experts do not know the source model of each image. We employ two metrics, including 1) subject similarity defined as the similarity between subjects from generated images and given images, and 2) prompt consistency defined as the consistency among prompts and generated images. Table 6 reports the average score for all the images. We can find that MoSLoRA outperforms LoRA on both metrics. In particular, MoSLoRA gets an average winning ratio of 45.1% on prompt consistency, which is 34.3% than LoRA. Please refer to Appendix E for the detailed prompts and corresponding generated images from LoRA and MoSLoRA.
@@ 5.2 LoRA and its Variants
The core of LoRA is to update the mergeable and low-rank branches to model the weight updates. Hu et al. (2022) initialize the branch as a product of two low-rank matrices. The following variants can be categorized into: 1) introducing training skills, such as setting different learning rates (Hayou et al., 2024) and adding random noise (Lin et al., 2024); 2) searching ranks, such as DyLoRA (Valipour et al., 2023) and AdaLoRA (Zhang et al., 2023); and 3) new designs for the branch, such as LoKr (Yeh et al., 2023), LoHa (Yeh et al., 2023), VeRA (Kopiczko et al., 2023) and (Liu et al., 2024). LoKr and LoHa employ Kronecker and Hadamard products to replace the vanilla matrix product, respectively. DoRA decomposes the pretrained weight into its magnitude and directional components and fine-tunes them separately.
In this paper, we decompose LoRA into subspaces via structural re-parameterization and design a learnable mixer to fuse information from more subspaces and more flexibly.
@@ A.2 Visual Instruction Tuning
  • SEED Bench (Li et al., 2023): 19K multiple choice questions with accurate human annotations, which spans 12 evaluation dimensions including the comprehension of both the image and video modality. In this paper, we use the image modality only and report the accuracy.
@@
  • SciQA (ScienceQA) (Lu et al., 2022): 21k multimodal multiple choice questions with diverse science topics and annotations of their answers with corresponding lectures and explanations. We report the accuracy of the test set.
@@ Appendix C Initialize Mixer as Zero Matrix
Appendix D Differences with FLoRA
FLoRA (Si et al., 2024) introduces an extra core based on Tucker decomposition to maintain the consistent topological structure with the original space. The core is quite similar to our mixer. The differences are:
  • motivation: FLoRA aims to maintain the structural integrity of the involved high-dimensional spaces, while we try to analyze LoRA and two-subspaces-mixing LoRA in the view of subspace.
  • initialization: FLoRA initializes the core as the zero matrix, while MoSLoRA employs the Kaiming uniform distribution and orthogonal matrix. As shown in Table 4, MoSLoRA outperforms FLoRA on the commonsense reasoning tasks.
Appendix E Cases of Generated Images

### v2

@@ 4.1 Commonsense Reasoning
We fine-tune LLaMA-3 8B instruction version model (AI@Meta, 2024) for the commonsense reasoning question answering tasks. We first train the model using 170k training samples (Hu et al., 2023), and then test the fine-tuned model on 8 commonsense reasoning question answering benchmarks (refer to Appendix A.1 for details). The 170k training set is the mixture of the training sets of these benchmarks. Besides LoRA (Hu et al., 2022), we also compare MoSLoRA with various baselines, including: 1) LoKr (Yeh et al., 2023) which employs Kronecker products for matrix decomposition of $\mathbf{A}\mathbf{B}$; 2) LoHa (Yeh et al., 2023) which decomposes the vanilla LoRA into the Hadamard product of two LoRA branches; 3) FLoRA (Si et al., 2024) which introduces an extra core based on Tucker decomposition to maintain the consistent topological structure with the original space 4) AdaLoRA (Zhang et al., 2023) which parameterizes the incremental updates of the pre-trained weight matrices in the form of singular value decomposition; and 5) DoRA (Liu et al., 2024) which decomposes the pretrained weight into its magnitude and directional components and fine-tunes both of them.
@@ 4.2 Visual Instruction Tuning
There are two stages in training LLaVA: 1) pretrain a two-layer MLP to project visual features to language space, and 2) optimize LLM and visual encoder (optional) for visual instruction tuning. In this paper, we employ the pretrained projector provided in XTuner (Contributors, 2023b), and then conduct visual instruction tuning on the LLM backbone and visual encoder, simultaneously. For the LLM backbones, we adapt the LLaMA3 8B (AI@Meta, 2024) and InternLM2 7B (Cai et al., 2024) using the off-the-shelf projecters 333pretrained projecters. For the visual encoder, we employ the ViT 444openai/clip-vit-large-patch14-336 (Dosovitskiy et al., 2021) large version. Due to limited resources, we fintune both the LLM backbone and visual encoder via LoRA/MoSLoRA on the 665K instruction-following data (Liu et al., 2023a), rather than optimize all the parameters in LLMs. For InternLM2, we employ the 4-bit QLoRA (Dettmers et al., 2023) and corresponding QMoSLoRA (QLoRA+MoSLoRA). Based on the results in Table 3, we initialize the mixer as the orthogonal matrix and Kaiming uniform distribution, separately. For specific hyperparameters, please refer to the Appendix B.2. It takes around 20 hours to fine-tune using 4 Nvidia A100 80G GPUs.
@@
After visual instruction tuning, we evaluate the trained model on 9 popular benchmarks, including MMBench EN/CN (Liu et al., 2023b), SEED Bench (Li et al., 2023a), AI2D (Kembhavi et al., 2016), SciQA (Lu et al., 2022b), TextVQA (Singh et al., 2019), MathVista testmini (Lu et al., 2023), MM-Vet (Yu et al., 2023), and MME (Fu et al., 2023). All the evaluations are done using the VLMEvalKit (Contributors, 2023a). Please refer to Appendix A.2 for the details of the dataset and the reported metrics. Specifically, we scale the MME scores to 100 to calculate the average score.
@@ 4.3 Subject-driven Generation
Figure 6 shows one case of a dog from the DreamBooth dataset (Ruiz et al., 2023). We finetune the SDXL555stable-diffusion-xl-base-1.0 model (Podell et al., 2023) via LoRA and MoSLoRA. In MoSLoRA, the mixer is initialized as an orthogonal matrix. During finetuning, the learning rate is 1e-4, and the batch size is 4. We train the model for 500 steps, which costs around 16 minutes using 1 80G A100 GPU. During generation, we infer 50 steps for the given prompts. Compared to vanilla LoRA, we can find that our proposed MoSLoRA captures more details of the subject and better conforms to the given prompt. Specifically, MoSLoRA learns more details about the given cat, including the color of the hairs around the neck and the shape of the paw. Meanwhile, the images from MoSLoRA are more consistent with the given prompts, such as the wet (thus clumped) hair and the floating gesture (spread hands).
@@
We also perform human evaluation on the generated images. First, we choose four subjects (i.e. cat, dog, grey sloth plushie, and can) from the DreamBooth dataset (Ruiz et al., 2023) and fine-tune the SDXL model, respectively. Then, we randomly select 8 prompts to generate the corresponding images. After that, 15 human experts are asked to independently score win/tie/loss for the paired images from LoRA and MoSLoRA. During evaluation, we shuffle these pairs and keep that these experts do not know the source model of each image. We employ two metrics, including 1) subject similarity defined as the similarity between subjects from generated images and given images, and 2) prompt consistency defined as the consistency among prompts and generated images. Table 6 reports the average score for all the images. We can find that MoSLoRA outperforms LoRA on both metrics. In particular, MoSLoRA gets an average winning ratio of 45.1% on prompt consistency, which is 34.3% than LoRA. Please refer to Appendix D for the detailed prompts and corresponding generated images from LoRA and MoSLoRA.
@@ 5.2 LoRA and its Variants
The core of LoRA is to update the mergeable and low-rank branches to model the weight updates. Hu et al. (2022) initialize the branch as a product of two low-rank matrices. The following variants can be categorized into: 1) introducing training skills, such as setting different learning rates (Hayou et al., 2024) and adding random noise (Lin et al., 2024); 2) searching ranks, such as DyLoRA (Valipour et al., 2023) and AdaLoRA (Zhang et al., 2023); and 3) new designs for the branch, such as LoKr (Yeh et al., 2023), LoHa (Yeh et al., 2023), VeRA (Kopiczko et al., 2023), DoRA (Liu et al., 2024), and FLoRA (Si et al., 2024). LoKr and LoHa employ Kronecker and Hadamard products to replace the vanilla matrix product, respectively. DoRA decomposes the pretrained weight into its magnitude and directional components and fine-tunes them separately.
We also notice a very recent concurrent work FLoRA (Si et al., 2024). Differences between MoSLoRA and FLoRA are as follows: 1) initialization methods and the corresponding motivation: The motivation of FLoRA is to maintain the structural integrity of the original high-dimensional parameter spaces (i.e., 2D Convolution) by introducing the core spaces. Differently, MoSLoRA is motivated by the observation of exploratory experiments in Section 2.2, where we find the two-subspaces-mixing LoRA in the view of more fine-grained subspace has the potential to learn complex features. As a result, we introduce an extra learnable mixer initialized as Kaiming uniform distribution or orthogonal matrix, which we empirically (Section 3.2) find vital for the final performance. 2) analysis and derivation: The design of our mixer comes from the process of analyzing and deriving the two-subspaces-mixing LoRA. We revisit the two-subspaces-mixing strategy in a more fine-grained (rank=1) view and composed view, and then we find that both LoRA and two-subspaces-mixing LoRA are equivalent to inserting fixed mixer matrices. Based on that, we propose MoSLoRA, which employs a learnable mixer to fuse more subspaces and more flexibly. 3) models and datasets: For visual-instruction tuning tasks, FLoRA fine-tunes LLaVA-1.5-7B (based on Vicuna-1.5-7B (Peng et al., 2023)) and evaluates on seven vision language benchmarks: VQAv2 (Goyal et al., 2019), GQA (Hudson and Manning, 2019), VisWiz (Gurari et al., 2018), SQA (Lu et al., 2022a), TextVQA (Singh et al., 2019), POPE (Li et al., 2023b), and MMBench (Liu et al., 2023b). Differently, We select LLaMA3 8B and InternLM2 as the language models and evaluate the fine-tuned models on MMBench EN/CN (Liu et al., 2023b), SEED Bench (Li et al., 2023a), AI2D (Kembhavi et al., 2016), SciQA (Lu et al., 2022b), TextVQA (Singh et al., 2019), MathVista testmini (Lu et al., 2023), MM-Vet (Yu et al., 2023), and MME (Fu et al., 2023). Besides, for language models, we evaluate MoSLoRA on LLaMA3 8B, while FLoRA focuses on DeBERTaV3 (He et al., 2023).
@@ A.2 Visual Instruction Tuning
  • SEED Bench (Li et al., 2023a): 19K multiple choice questions with accurate human annotations, which spans 12 evaluation dimensions including the comprehension of both the image and video modality. In this paper, we use the image modality only and report the accuracy.
@@
  • SciQA (ScienceQA) (Lu et al., 2022b): 21k multimodal multiple choice questions with diverse science topics and annotations of their answers with corresponding lectures and explanations. We report the accuracy of the test set.
@@ Appendix C Initialize Mixer as Zero Matrix
Appendix D Cases of Generated Images